OpenStudy (anonymous):

I am 12 years old. can someone please help me solve: 1 11/15 divided by 1 19/20? The answer book says 8/9 but I can't seem to get the answer on my own. Thanks. but I want to see the

OpenStudy (anonymous):

1 11/15 = 26/15 1 19/20 = 39/20 do you see how i got that?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

okay so now, 26/15 divided 39/20 you know that when dividing by fractions, you multiply by the reciprocal of the second fraction so 26/15 divided 29/20 is equivalent to 26/15 x 20/39

OpenStudy (anonymous):

first you have to turn both fractions into improper fractions. 1 11/15= 26/15, 1 19/20 =39/20. To divide them, it is the same as taking the first fraction and multiplying that by the reciprocal of the second one, so it is 26/15 x 20/39 and you multiply that straight across, then simplify.

OpenStudy (anonymous):

^yup lol

OpenStudy (anonymous):

i multiplied to get 104/117 and can't simplify.